# A Survey on Adversarial Contention Resolution

### v3

@@ 5. Conclusion
We have surveyed the literature on adversarial contention resolution by delineating results based on key model assumptions, metrics, and algorithm-design choices. The problem of contention resolution remains an active topic of research, and here we briefly discuss some potential directions for future research.
Machine learning (ML) may have an important role to play in the design of contention-resolution algorithms; certainly, the use of ML is being explored in the development of future wireless standards, such as 6G (kato:ten, ). As discussed in Section 3.3.2, Gilbert et al. (gilbert2021contention, ) demonstrate the impact of system-size predictions, which might be offered by a black-box ML module. A natural question is whether there are other useful predictions that could give improved performance, or at least performance that is parameterized by the accuracy of the predictions. This has been done recently for denial-of-service attacks in the traditional client-server model (chakraborty:bankrupting, ). Can predictions about whether a slot will be noisy (or jammed) yield improved robustness?
@@
In many wireless systems, collisions have significant cost. For example, in WiFi networks, the entire packet must be sent before the sender learns of success or failure (due to a collision). Under coded networks, multiple packets may be sent simultaneously; therefore, a collision ruins what could otherwise be multiple successful transmissions in a slot. In mobile settings, failure to transmit a packet may mean incurring a significant delay before the sender is next within range of the intended receiver. Therefore, modeling collision cost is another angle that presents some interesting open problems in the area of contention resolution.
@@
Adopting different channel models can yield distinct upper and lower bounds compared to the classical model, as we have seen in the case of the SINR model (e.g., (fineman:contention, ; jurdzinski2015cost, )). The power sensitive model explored by De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) and by Tsybakov tsybakov:multiplicity) is another example of a interesting channel model. In the latter case, exploring via experiment the degree to which this model holds in practice may be interesting, as it seems to impart a non-trivial amount of power (e.g., we can count the number of users in $O(1)$ slots rather than ${\texttt{poly}({\log n})}$ slots). Even coarser-grained feedback —rather an exact value—for the number of senders could be useful. For example, in the classical channel, it is hard to tell if noise comes from many users competing for access or from jamming; however, (even rough) estimates could help clarify this.
Another dimension to consider is traffic priority. Consider a packet that contains voice-over-IP data, and another packet holding old data being backed up on a remote server. Which packet should get access to the shared channel first? Intuitively, the first packet is time sensitive and should have priority; if it is delayed too long, the packet contents are useless. This prioritization will become increasingly important to wireless networks, as a combination of limited bandwidth and a growing number of devices forces a differentiated treatment of traffic. While this aspect has been explored in (DBLP:conf/spaa/AgrawalBFGY20, ), there are several outstanding questions regarding robustness to adversarial jamming, heterogeneous packet sizes, and potential trade-offs with the amount of delay packets may suffer as a result.
Energy efficiency is likely to remain an important consideration in wireless systems. The exact limits of trading off between robustness to worst-case noise, energy efficiency, and latency still remain unknown. Can we obtain $o(\log n)$ number of channel-access attempts per packet despite jamming, while retaining asymptotically optimal latency? Related to this aspect, in some wireless standards, transmission power can be increased (and decreased), allowing devices to compensate for signal attenuation. This capability might be useful in overcoming jamming attacks, where good participants incur higher energy costs to transmit at higher power in order to overcome malicious noise. Is it possible to design algorithms in the spirit of (ChangJP19, ; BenderFGY19, ), where the sending costs are grow very slowly with the jamming costs? Alternatively, can the latency be asymptotically improved over existing jamming-resistant results?

### v4

@@ 5. Conclusion
We have surveyed the literature on adversarial contention resolution by delineating results based on key model assumptions, metrics, and algorithm-design choices. The problem of contention resolution remains an active topic of research, and here we briefly discuss some potential directions for future work.
Machine learning (ML) may have an important role to play in the design of contention-resolution algorithms; certainly, the use of ML is being explored in the development of future wireless standards, such as 6G (kato:ten, ). As discussed in Section 3.3.2, Gilbert et al. (gilbert2021contention, ) demonstrate the impact of system-size predictions, which might be offered by a black-box ML component. A natural question is whether there are other useful predictions that can give improved performance, or at least performance that is parameterized by the accuracy of the predictions. This has been done recently for denial-of-service attacks in the traditional client-server model (chakraborty:bankrupting, ). Can predictions about whether a slot will be noisy (or jammed) yield improved robustness?
@@
In many wireless systems, collisions have significant cost. For example, in WiFi networks, the entire packet must be sent before the sender learns of success or failure (due to a collision). For instance, in mobile settings, failure to transmit a packet may mean incurring a significant delay before the sender is next within range of the intended receiver. Therefore, modeling and minimizing collision cost presents some interesting open problems.
@@
Adopting different channel models can yield distinct upper and lower bounds compared to the classical model, as we have seen in the case of the SINR model (e.g., (fineman:contention, ; jurdzinski2015cost, )). The power-sensitive model explored by De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) and by Tsybakov (tsybakov:multiplicity, ) is another example of an interesting channel model. In the latter case, exploring via experiment the degree to which this model holds in practice may be interesting, as it seems to impart a non-trivial amount of power (e.g., we can count the number of users in $O(1)$ slots rather than ${\texttt{poly}({\log n})}$ slots). Even coarser-grained feedback —rather than an exact value—for the number of senders could be useful. For example, in the classical channel, it is hard to tell if noise comes from many users competing for access or from jamming; however, (even rough) estimates could help clarify this.
Another dimension to consider is traffic priority. Consider a packet that contains voice-over-IP data, and another packet holding old data being backed up on a remote server. Which packet should get access to the shared channel first? Intuitively, the first packet is time sensitive and should have priority; if it is delayed too long, the packet contents are useless. This prioritization may become increasingly important to wireless networks, as a combination of limited bandwidth and a growing number of devices forces a differentiated treatment of traffic. While this aspect has been explored by Agrawal et al. (DBLP:conf/spaa/AgrawalBFGY20, ), there are several outstanding questions regarding robustness to adversarial jamming, heterogeneous packet sizes, and potential trade-offs with the amount of delay packets may suffer as a result.
Energy efficiency is likely to remain an important consideration in wireless systems. The exact limits of trading off between robustness to worst-case noise, energy efficiency, and latency still remain unknown. Can we obtain $o(\log n)$ number of channel-access attempts per packet despite jamming, while retaining asymptotically optimal latency? Related to this aspect, in some wireless standards, transmission power can be adjusted, allowing devices to compensate for signal attenuation. This capability might be useful in overcoming jamming attacks, where good participants incur higher energy costs to transmit at higher power in order to overcome malicious noise. Is it possible to design algorithms in the spirit of (ChangJP19, ; BenderFGY19, ), where the sending costs are grow very slowly with the jamming costs? Alternatively, can the latency be asymptotically improved over existing jamming-resistant results?
Acknowledgements.
The authors are grateful to the anonymous reviewers for their feedback.